A 300. mg sample of caffeine was dissolved in 10.0 g of camphor (Kf = 39.7 °C/m), decreasing the freezing point of camphor by 3.07 °C. What is the molar mass of caffeine?

1
Expert's answer
2022-04-01T13:11:03-0400

ΔT = 3.07

Kf = 39.7

m – molality

moles = 0.300 / M

mass of solvent = 10.0 g = 0.01 kg

m = 0.300 / (M * 0.01)

ΔT = Kf * m

3.07 = (39.7 * 0.300) / (M * 0.01)

M = 387.9